OpenStudy (anonymous):

Ok. A determinants question(Link below). What I have done? Well, I tried Row1 = Row 1+Row 2+row 3 but nothing great. Can you guys guide me to get the answer??

OpenStudy (anonymous):

http://mypat.in/PaperImages/3/425/Q_4_004.gif

OpenStudy (unklerhaukus):

can you factor all of the elements

OpenStudy (anonymous):

@UnkleRhaukus ,how exactly??

OpenStudy (unklerhaukus):

\[\Delta=\left|\begin{array}\\1+a^2+a^4&1+ab+a^2b^2&1+ac+a^2c^2\\1+ab+a^2b^2&1+b^2+b^4&1+bc+b^2c^2\\1+ac+a^2c^2&1+bc+b^2c^2&1+c^2+c^4\end{array}\right|\]

OpenStudy (unklerhaukus):

yeah maybe not, i really dont want to do it the normal way

OpenStudy (anonymous):

Normal way-->never :D It would take ages :P

OpenStudy (unklerhaukus):

but it would work in theory, but i would make about ten mistakes

OpenStudy (anonymous):

I am sure you play around with determinant properties like Row1 = Row 1 + Row 2+Row3 . I mean some tweaking of that sort

OpenStudy (anonymous):

@apoorvk

OpenStudy (inkyvoyd):

can we just do it the foolish way?

OpenStudy (apoorvk):

Hmm, tried C1 - C2?

OpenStudy (anonymous):

@inkyvoyd Sorry, please, In my exam I have only 3 min(max) to solve this question @apoorvk ,let me give it a try and get back to you

OpenStudy (anonymous):

@apoorvk , I did that. But nothing can be taken common out :(

OpenStudy (inkyvoyd):

I'm going to try this problem

OpenStudy (inkyvoyd):

Have no fear, for inky is here!

OpenStudy (apoorvk):

I got common stuff out through C1-C2 and C2-C3. but, i am still stuck with a trinomial in each ditch..

OpenStudy (inkyvoyd):

@shivam_bhalla I HAVE AN IDEA

OpenStudy (inkyvoyd):

aei+bfg+cdh-ceg-bdi-afh =a(ei-fh)+b(fg-dj)+c(dg-eg)

OpenStudy (inkyvoyd):

b=d f=h

OpenStudy (inkyvoyd):

c=g

OpenStudy (inkyvoyd):

aei+bfg+cdh-ceg-bdi-afh aei+bf^2+cbf-ec^2-ah^2

OpenStudy (anonymous):

@inkyvoyd , what is a,b,c,d,e,f,g?? Are you expanding the determinant??

OpenStudy (inkyvoyd):

I'm wikipediaing, and not making any progress.

OpenStudy (anonymous):

LOL. @ "wikipediaing" . :)

OpenStudy (apoorvk):

wolfram!!!

OpenStudy (anonymous):

How to wolfram this ??

OpenStudy (anonymous):

@radar sir, any clue ??

OpenStudy (experimentx):

factor out the commons!!!

OpenStudy (anonymous):

@experimentX , how exactly ?? I mean row wise or column wise. And moreover, what to factor out ??

OpenStudy (experimentx):

lol ... factoring out would lead to 3 more matrices ..thats ugly... must be some pattern!!

OpenStudy (anonymous):

Oh. Now get what you meant by factoring .:P I tried that too and observed what you said :)

OpenStudy (experimentx):

it seems transpose of itself!!

OpenStudy (apoorvk):

 this is how far i got then..